Holding constant the cost of the bundle in each country, compute the real exchange rates that would result from the two nominal exchange rates in the following table.

3. Computing real exchange rates 


Consider a bundle of consumer goods that costs $90 in the United States. The same bundle of goods costs CNY 105 in China.


 Holding constant the cost of the bundle in each country, compute the real exchange rates that would result from the two nominal exchange rates in the following table. 


  Cost of Bundle in U.S.   Cost of Bundle in China   Nominal Exchang             Real Exchange Rate

       (Dollars)              (Yuan)            (Yuan per dollar)  (Bundles of Chinese goods per bundle of U.S. goods)

          90                   105                   7.00

          90                   105                   10.50
Answer #2

Cost of basket in US at 7 nominal rate = Cost of basket in US dollar × Yuan per dollar

                                                              = 90 × 7

                                                              = 630

Real exchange rate = (Cost in US) / (Cost in C) = 630 / 105 = 6 (Answer)

Cost of basket in C at 10.50 nominal rate = Cost of basket in US dollar × Yuan per dollar

                                                                   = 90 × 10.50

                                                                   = 945

Real exchange rate = (Cost in US) / (Cost in C) = 945 / 105 = 9 (Answer)
Answer #1

Real exchange rate = nominal exchange rate *cost in US / cost in China

1) real exchange rate = 7*90/105= 6

2) real exchange rate = 10.5*90/105= 9